OpenStudy (anonymous):

started a new question so I could give you another medal for checking my answers :]

OpenStudy (anonymous):

@jeweaver6895 Answers: tan(a)=x/y sin(b)=y/z cos(b)=x/z tan(b)=y/x

OpenStudy (anonymous):

Perfect! Great job, you've got a good handle on this! :)

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

Thank you!! You too Kstate :] Thanks again for teaching me how to do this. I've been stuck on this for ages and even my teacher wasn't much help! Have a great day.

OpenStudy (anonymous):

Np

OpenStudy (anonymous):

You bet! Glad I was able to help! Feel free to tag me in any future questions! Have a great day as well!

OpenStudy (anonymous):

Thanks so much! :] Bye!